CRY+VIS Interface Meeting
Date: 2022/4/8 13:00-13:40
Participants: Takahashi, Ikeda, Sato, Yokozawa, Node, Aso, Ushiba, Hirata, Uchiyama, Tomaru, Akutsu, Miyoki
- IY: Magnet for the TM actuator was reglued. AA for the geophone signal was replaced. Go to finalize (Baffle, Frame adjustment).
- EY: TM actuator is under investigation.
- Type-B/Bp/C: Feedthrough flange for MCO was replaced. MCO suspension was realigned. Power supply for PRM/PR3 was replaced. BFH LVDT is not working. TF for SR2 was checked.
- Replacement of flip cables: Feedthrough flange was replaced in IFI. Cables for TMS were ordered.
- TM height setting: ITMY.
- Oplev: Calibration in EY. Windshield in the next.
- Commissioning in air:
- Commissioning in vacuum:
- Acceptance check:
- PF actuator: At the in-vacuum commissioning.
- New vacuum pumps for V-chambers: Opening of the GV in IXV/EXV.
- Heater control: Test(IXV/EXV).
- TMS: Pressure was increased in EXT. Monitor is not working now.
Date: 2022/4/8 14:00-14:50
Participants: Takahashi, Uchiyama, Fabian, Terrence, Aso, Tanaka, Washimi, Luccia
- Closed the ICF 203 flanges on the corrugate tube in IY.
- Checked the transfer functions in SR2.
- Designing the limit switch jig for the traverse.
- BFH LVDT is not working in PRM.
- Replaced the feedthrough flange in MCF.
- Replaced the feedthrough flange in IFI.
- Inspected the MCO suspension.
- Type-A team: Finalization of IY.
- Type-B team: BFH LVDT in PRM. SR3 FR.
Active isolation reserarch JGW-G2214020
(Date in Kamioka)
The week from 4/11
- Takahashi: 4/12-15
- Sato: 4/12-15
- Hirata: 4/12-15
The week from 4/18
- Takahashi: 4/19-22
- Sato: 4/19-22
- Hirata: 4/19-22